# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v4

@@ Abstract
Membership inference attacks (MIAs) are used to test practical privacy of machine learning models. MIAs complement formal guarantees from differential privacy (DP) under a more realistic adversary model. We analyse MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning. We show that the vulnerability of non-DP models when measured as the attacker advantage at fixed false positive rate reduces according to a simple power law as the number of examples per class increases, even for the most vulnerable points, but the dataset size needed for adequate protection of the most vulnerable points is very large.
@@ 1 Introduction
Membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and differential privacy (DP; Dwork et al., 2006) provide complementary means of deriving lower and upper bounds for the privacy loss of a machine learning algorithm. Yet, the two operate under slightly different threat models. DP implicitly assumes a very powerful adversary with access to all training data except the target point and provides guarantees against every target point. MIAs assume an often more realistic adversary model with access to just the data distribution, and the unknown training data becoming latent variables introducing stochasticity into the attack, but the practical evaluation is statistical and cannot provide universal guarantees.
In this paper, we seek to explore the area between DP and typical MIA evaluations. Inspired by an empirical finding that average MIA vulnerability of neural network fine-tuning strongly reduces as the number of samples in the target class increases (see Figure 1), we develop theory of optimal MIA against a simple model of neural network fine-tuning and reproduce the decrease in vulnerability. Furthermore, the theoretical model predicts that the vulnerability of all individual samples should reduce as the number of samples increases, which we are able to verify empirically.
Figure 1: We observe a power-law relation between MIA vulnerability and $S$ (shots, i.e., examples per class) when attacking a ViT-B Head fine-tuned using LiRA. The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over six seeds.
@@
We focus on transfer learning using fine-tuning because this is increasingly used for all practical applications of deep learning and especially important when labeled examples are limited, which is often the case in privacy-sensitive applications. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep transfer learning models to MIA. In particular, we theoretically study the relationship between the number of examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning, and derive a power-law relationship (Figure 1) in the form
@@
Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020). A larger generalisation error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018). Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., ; Bagdasaryan et al., 2019). Feldman and Zhang (2020) showed that neural networks trained from scratch can memorise a large fraction of their training data, while the memorisation is greatly reduced for fine-tuning. Additionally, Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots. Yu et al. (2023) studied the relationship between the MIA vulnerability and individual privacy parameters for different classes. Recently, worst-case MIA vulnerability has gained more attention (Guépin et al., 2024; Meeus et al., 2024; Azize and Basu, 2024). Nonetheless, the prior works do not consider the rate of change in the vulnerability evaluated at a low fpr, as dataset properties change. Our work significantly expands on these works by explicitly identifying a quantitative relationship between dataset properties and MIA vulnerability, i.e., the power-law in Equation 1.
List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art score-based MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), which are a strong realistic threat model. We first analytically derive the power-law relationship in Equation 1 for both MIAs by introducing a simplified model of the optimal membership inference (Section 3). We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models by focusing on a transfer-learning setting for image classification task, where a large pre-trained neural network is fine-tuned on a sensitive dataset.
@@
    Power-law in a simplified model of the optimal MIA: We formulate a simplified model of MIA to quantitatively relate dataset properties and MIA vulnerability. In this model LiRA is the optimal attack. For this model, we prove a power-law relationship for both average and worst-case between the LiRA as well as RMIA vulnerability and the number of examples per class (See Section 3.4).
@@
    MIA experiments on the average case vulnerability: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power-law between the number of examples per class and the vulnerability to MIA (see Figures 1 and 4.2). We fit a regression model which follows the functional form of the theoretically derived power-law. We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Section 4.3).
@@
    MIA experiments on the worst-case vulnerability: We extend the experiments to worst-case individual sample vulnerabilities and observe a similar decrease in vulnerability for quantiles of vulnerable data points and a slower decrease for the maximum individual vulnerability (Section 4.4). By extrapolation we find that an adequate protection of the most vulnerable samples would require an extremely large dataset (Section 4.5).
@@ 2 Background
Notation for the properties of the training dataset $\mathcal{\mathcal{D}}$: 1 $C$for the number of classes 2 $S$for shots (examples per class) 3 $\mathcal{|\mathcal{D}|}$for training dataset size ($\mathcal{|\mathcal{D}|}=CS$). We denote the number of MIA shadow models with $M$.
@@ 3 Theoretical analysis
In this section, we seek to theoretically understand the impact of the dataset properties on the MIA vulnerability. It is known that different data points exhibit different levels of MIA vulnerability depending on the underlying distribution (e.g.  Aerni et al., 2024; Leemann et al., 2024). Therefore, we start with analysing per-example MIA vulnerabilities. In order to quantitatively relate dataset properties to these vulnerabilities, a simplified model is formulated. Within this model, we prove a power-law between the per-example vulnerability and the number $S$ of examples per class. Finally, the per-example power-law is analytically extended to average-case MIA vulnerability, for which we provide empirical evidence in Section 4. We primarily focus on the analysis of LiRA, since it is the optimal attack in our simplified model. In Appendix B we show that similar theoretical results also hold for RMIA.
@@ 3.1 Preliminaries
First, let us restate the MIA score from LiRA as defined by Carlini et al. (2022). Denoting the logit of a target model $\mathcal{M}$ applied on a target data point $(\bm{x},y)$ as $\ell(\mathcal{M}(\bm{x}),y)$, LiRA computes the MIA score as the likelihood ratio
@@ 3.2 Computing the tpr for LiRA
Using the LiRA formulation of Equation 4, the tpr for the target point $(\bm{x},y)$ for LiRA is defined as
@@
where $\tau$ is a threshold that defines a rejection region of the likelihood ratio test, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Section A.1 for derivation).
@@ 3.3 Per-example MIA vulnerability
where $Z$ has the standard location and unit scale, and $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and $\sigma_{\mathrm{in}}(\bm{x}),\sigma_{\mathrm{out}}(\bm{x})$ are the locations and scales of IN/OUT distributions of $t_{\bm{x}}$. We assume that the target and shadow datasets have a sufficient number of examples. This allows us to also assume that $\hat{\sigma}(\bm{x})=\hat{\sigma}_{\mathrm{in}}(\bm{x})=\hat{\sigma}_{\mathrm{out}}(\bm{x})$ and $\sigma(\bm{x})=\sigma_{\mathrm{in}}(\bm{x})=\sigma_{\mathrm{out}}(\bm{x})$, where $\hat{\sigma}(\bm{x})$ is the standard deviation of $t_{\bm{x}}$ estimated from shadow models and $\sigma(\bm{x})$ is the true scale parameter of $t_{\bm{x}}$. (See Section A.2 for the validity of these assumptions). The following result reduces the LiRA vulnerability to the location and scale parameters of $t_{\bm{x}}$.
@@
See Section A.3. ∎
Here we assume that an attacker trains shadow models with the true underlying distribution. However, in real-world settings the precise underlying distribution may not be available for an attacker. We relax this assumption in Section A.4 so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma 1.
@@ 3.4 A simplified model of the optimal membership inference
This simplified model resembles a linear (Head) classifier often used in transfer learning when adapting to a new dataset. We also focus on the linear (Head) classifier in our empirical evaluation in Section 4. In the linear classifier, we find a matrix $W$ and biases $b$, to optimize the cross-entropy between the labels and logits $Wv+b$, where $v$ denotes the feature space representation of the data. In the simplified model, the rows of $W$ are replaced by the cluster means and we do not include the bias term in the classification.
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (9)
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\leq-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{||\bm{x}-m_{\bm{x}}||}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\\ +\log(1+\xi(S)).$ (11)
This implies that if $||\bm{x}-m_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded. Hence we can significantly reduce the MIA vulnerability of all examples in this non-DP setting by simply increasing the number of examples per class.
Now the following corollary extends the power-law to the average-case MIA vulnerability. We will also empirically validate this result in Section 4.
@@ 4 Empirical evaluation of MIA vulnerability and dataset properties
We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream datasets and assess the MIA vulnerability using LiRA and RMIA with $M=256$ shadow/reference models. We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $>80\%$ (see Table A2).
@@ 4.1 Experimental setup
We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015). The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section D.1. We optimise the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section C.2). We provide the the code for reproducing the experiments in the supplementary material.
4.2 Experimental results
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section 2) using LiRA. We make the following observations:
  • A larger number of $S$ (shots) decrease the vulnerability in a power law relation as demonstrated in Figure 1. We provide tabular data and experiments using ResNet-50 in the Appendix (Figures A.1, A3 and A4).
  • Contrary, a larger number of $C$ (classes) increases the vulnerability as demonstrated in Figure 2 with tabular data and experiments using ResNet-50 in the Appendix (Figures A.2, A5 and A6). However, the trend w.r.t. $C$ is not as clear as with $S$.
RMIA In Figure 3 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$. We observe the power-law for both attacks, but the RMIA is more unstable than LiRA (especially for lower fpr). More results for RMIA are in Figures A.5, A.6 and A.7 in the Appendix.

### v5

@@ Abstract
Membership inference attacks (MIAs) are used to test practical privacy of machine learning models. MIAs complement formal guarantees from differential privacy (DP) under a more realistic adversary model. We analyse MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning. We show that the vulnerability of non-DP models when measured as the attacker advantage at a fixed false positive rate reduces according to a simple power law as the number of examples per class increases. A similar power-law applies even for the most vulnerable points, but the dataset size needed for adequate protection of the most vulnerable points is very large.
@@ 1 Introduction
Figure 1: We observe a power-law relation between MIA vulnerability and examples per class (denoted as $S$ or shots) when attacking a fine-tuned ViT-B Head using LiRA. Each colored line denotes a different fine-tuning dataset where $C$ specifies the number of classes. The solid line is median and the error bars the min/max bounds for the Clopper-Pearson CIs over six seeds.
Membership inference attacks (MIAs; Shokri et al., 2017; Carlini et al., 2022) and differential privacy (DP; Dwork et al., 2006) provide complementary means of deriving lower and upper bounds for the privacy loss of a machine learning algorithm. Yet, the two operate under slightly different threat models. DP implicitly assumes a very powerful adversary with access to all training data except the target point and provides guarantees against every target point.
MIAs assume an often more realistic adversary model with access to just the data distribution and the unknown training data becoming latent variables that introduce stochasticity into the attack. However, the practical evaluation is statistical and cannot provide universal guarantees.
In this paper, we seek to explore MIA vulnerability to extrapolate this gap. Inspired by an empirical finding that average MIA vulnerability of neural network fine-tuning strongly reduces as the number of samples in the target class increases (see Figure˜1), we develop theory of optimal MIA against a simple model of neural network fine-tuning and reproduce the decrease in vulnerability. Furthermore, the theoretical model predicts that the vulnerability of all individual samples should reduce as the number of samples increases, which we are able to verify empirically.
@@
We focus on transfer learning using fine-tuning because this is increasingly used for all practical applications of deep learning and especially important when labeled examples are limited, which is often the case in privacy-sensitive applications. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep transfer learning models to MIA. In particular, we theoretically study the relationship between the number of examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning, and derive a power-law relationship (Figure˜1) in the form
@@
Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2025), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020). A larger generalisation error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018). Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., 2021; Bagdasaryan et al., 2019). Feldman and Zhang (2020) showed that neural networks trained from scratch are required to memorize a significant fraction of their training data to obtain high utility, while the memorisation is greatly reduced for fine-tuning. Additionally, Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots. Yu et al. (2023) studied the relationship between the MIA vulnerability and individual privacy parameters for different classes. Recently, worst-case MIA vulnerability has gained more attention (Guépin et al., 2024; Meeus et al., 2024; Azize and Basu, 2024). Nonetheless, the prior works do not consider the rate of change in the vulnerability evaluated at a low fpr, as dataset properties change. Our work significantly expands on these works by a) explicitly identifying a quantitative relationship between dataset properties and MIA vulnerability, i.e., the power-law in Equation˜1, and b) focusing on the worst-case vulnerability, both evaluated at a low FPR. This in turn allows us to extrapolate MIA vulnerability to DP guarantee.
List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art score-based MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), which are a strong realistic threat model. We first analytically derive the power-law relationship in Equation˜1 for both MIAs by introducing a simplified model of the optimal membership inference (Section˜3). We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models by focusing on a transfer-learning setting for image classification task, where a large pre-trained neural network is fine-tuned on a sensitive dataset.
@@
    Power-law in a simplified model of the optimal MIA: We formulate a simplified model of MIA to quantitatively relate dataset properties and MIA vulnerability. In this model LiRA is the optimal attack. For this model, we prove a power-law relationship for both average and worst-case between the LiRA as well as RMIA vulnerability and the number of examples per class (See Section˜3.4).
@@
    MIA experiments on the average case vulnerability: We conduct a comprehensive study of MIA vulnerability (tpr at a fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power-law between the number of examples per class and the vulnerability to MIA (see Figures˜1 and 4.2). We fit a regression model which follows the functional form of the theoretically derived power-law. We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Section˜4.3).
@@
    MIA experiments on the worst-case vulnerability: We extend the experiments to worst-case individual sample vulnerabilities and observe a similar decrease in vulnerability for quantiles of vulnerable data points and a slower decrease for the maximum individual vulnerability (Section˜4.4). By extrapolation we find that an adequate protection of the most vulnerable samples would require an extremely large dataset (Section˜4.5).
@@ 2 Background
Notation for the properties of the training dataset $\mathcal{\mathcal{D}}$: 1 CCfor the number of classes 2 SSfor shots (examples per class) 3 $\mathcal{|\mathcal{D}|}$for training dataset size ($\mathcal{|\mathcal{D}|}=CS$). We denote the number of MIA shadow models with $M$.
@@ 3 Theoretical analysis
In this section, we seek to theoretically understand the impact of the dataset properties on the MIA vulnerability. It is known that different data points exhibit different levels of MIA vulnerability depending on the underlying distribution (e.g. Aerni et al., 2024; Leemann et al., 2024). Therefore, we start with analysing per-example MIA vulnerabilities. In order to quantitatively relate dataset properties to these vulnerabilities, a simplified model is formulated. Within this model, we prove a power-law between the per-example vulnerability and the number $S$ of examples per class. Finally, the per-example power-law is analytically extended to average-case MIA vulnerability, for which we provide empirical evidence in Section˜4. We primarily focus on the analysis of (online) LiRA, since it is the optimal attack in our simplified model. We show that similar theoretical results also hold for RMIA in Appendix˜B and offline LiRA Section˜A.4.
@@ 3.1 Preliminaries
First, let us restate the MIA score from (online) LiRA as defined by Carlini et al. (2022). Denoting the logit of a target model $\mathcal{M}$ applied on a target data point $(\bm{x},y)$ as $\ell(\mathcal{M}(\bm{x}),y)$, LiRA computes the MIA score as the likelihood ratio
@@ 3.2 Computing the tpr for LiRA
Using the LiRA formulation of Equation˜4, the tpr for the target point $(\bm{x},y)$ for LiRA is defined as
@@
where $\tau$ is a threshold that defines a rejection region of the likelihood ratio test, and $\phi^{M}$ denotes the randomness in shadow set sampling and shadow model training (see Section˜A.1 for derivation).
@@ 3.3 Per-example MIA vulnerability
where $Z$ has the standard location and unit scale, and $\mu_{\mathrm{in}}(\bm{x}),\mu_{\mathrm{out}}(\bm{x})$ and $\sigma_{\mathrm{in}}(\bm{x}),\sigma_{\mathrm{out}}(\bm{x})$ are the locations and scales of IN/OUT distributions of $t_{\bm{x}}$. We assume that the target and shadow datasets have a sufficient number of examples. This allows us to also assume that $\hat{\sigma}(\bm{x})=\hat{\sigma}_{\mathrm{in}}(\bm{x})=\hat{\sigma}_{\mathrm{out}}(\bm{x})$ and $\sigma(\bm{x})=\sigma_{\mathrm{in}}(\bm{x})=\sigma_{\mathrm{out}}(\bm{x})$, where $\hat{\sigma}(\bm{x})$ is the standard deviation of $t_{\bm{x}}$ estimated from shadow models and $\sigma(\bm{x})$ is the true scale parameter of $t_{\bm{x}}$. (See Section˜A.2 for the validity of these assumptions). The following result reduces the LiRA vulnerability to the location and scale parameters of $t_{\bm{x}}$.
@@
See Section˜A.3. ∎
Here we assume that an attacker trains shadow models with the true underlying distribution. However, in real-world settings the precise underlying distribution may not be available for an attacker. We relax this assumption in Section˜A.5 so that the attacker only needs an approximated underlying distribution for the optimal LiRA as in Lemma˜1.
@@ 3.4 A simplified model of the optimal membership inference
This simplified model resembles a linear (Head) classifier often used in transfer learning when adapting to a new dataset. We also focus on the linear (Head) classifier in our empirical evaluation in Section˜4. In the linear classifier, we find a matrix $W$ and biases $b$, to optimize the cross-entropy between the labels and logits $Wv+b$, where $v$ denotes the feature space representation of the data. In the simplified model, the rows of $W$ are replaced by the cluster means and we do not include the bias term in the classification.
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (9)
@@
[…]
This implies that if $\bm{x}-m_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded. Hence we can significantly reduce the MIA vulnerability of all examples in this non-DP setting by simply increasing the number of examples per class.
Remark 3.
By Lemma˜1 and the proof of Theorem˜2, a necessary condition for the power-law is that $(\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x}))/\sigma(\bf{x})$ converges to zero at rate $O(1/S^{\alpha})$ with $\alpha>0$. In our simplified model, this holds with $\alpha=1/2$. However, $\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x})\rightarrow 0$ would not always be the case for larger neural networks trained from scratch. Therefore, we do not expect similar power-laws in more general training algorithms.
Now the following corollary extends the power-law to the average-case MIA vulnerability. We will also empirically validate this result in Section˜4.
@@ 4.1 Experimental setup
We focus on a image classification setting where we fine-tune pre-trained models on sensitive downstream datasets and assess the MIA vulnerability using LiRA and RMIA with $M=256$ shadow/reference models. We base our experiments on a subset of the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a test classification accuracy $80\%$ (see Table˜A2).
We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015). The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section˜D.1. We optimise the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section˜C.2). We provide the the code for reproducing the experiments in an open repository [^1] https://github.com/DPBayes/impact-dataset-properties-MI-vulnerability-deep-TL.
4.2 Experimental results
@@
Using the setting described above, we study how the number of classes and the number of shots affect the vulnerability (tpr at fpr as described in Section˜2) using LiRA. We make the following observations:
@@
[…]
RMIA In Figure˜3 we compare the vulnerability of the models to LiRA and RMIA as a function of the number of $S$ (shots) at $\textsc{fpr}=0.1$. We observe the power-law for both attacks, but the RMIA is more unstable than LiRA (especially for lower fpr). More results for RMIA are in Figures˜A.5, A.6 and A.7 in the Appendix.
Figure 3: LiRA and RMIA vulnerability ($(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$) as a function of shots ($S$) when attacking a ViT-B Head fine-tuned on different datasets. For better visibility, we split the datasets into two panels. We observe the power-law for both attacks, but the RMIA is more unstable than LiRA. The lines display the median over six seeds.